40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 197 of 355  Not logged in ELOG logo
ID Date Author Type Category Subject
  7949   Mon Jan 28 21:32:38 2013 jamieUpdateAlignmenttweaking of alignment into half PRC

[Koji, Jamie]

We tweaked up the alignment of the half PRC a bit.  Koji started by looking at the REFL and POP DC powers as a function of TT2 and PRM alignment. 
He found that the reflected beam for good PRC transmission was not well overlapped at REFL.  When the beam was well overlapped at REFL, there was clipping in the REFL path on the AS table.

We started by getting good overlap at REFL, and then went to the AS table to tweak up all the beams on the REFL pds and cameras.
This made the unlocked REFL DC about 40 count. This was about 10mV (=0.2mA) at the REFL55 PD.
This amazed Koji since we found the REFL DC (of the day) of 160 as the maximum of the day for a particular combination of the PRM Pitch and TT2 Pitch. So something wrong could be somewhere.

We then moved to the ITMX table where we cleaned up the POP path.  We noticed that the lens in the POP path is a little slow, so the beam is too big on the POP PD and on the POP camera (and on the camera pick-off mirror as well)
We moved the currently unused POP55 and POP22/110 RFPDs out of the way so we could move the POP RF PD and camera back closer to the focus.  Things are better, but we still need to get a better focus, particularly on the POP PD.

We found two irides on the oplev path. They are too big and one of these is too close to the POP beam. Since it does not make sense too to have two irides in vicinity, we pulled out that one from the post.

Other things we noticed:

  • The POP beam is definitely clipping in the vacuum, looks like on two sides.
  • We can probably get better layout on the POP table, so we're not hitting mirrors at oblique angles and can get beams on grid paths.

After the alignment work on the tables, we started locking the cavity. We already saw the improvement of the POPDC power from 1000 cnt to 2500 cnt without any realignment.
Once PRM is tweaked a little (0.01ish for pitch and yaw), the maximum POPDC of 6000 was achieved. But still the POP camera shows non-gaussian shape of the beam and the Faraday camera shows bright
scattering of the beam. It seems that the scattering at the Faraday is not from the main beam but the halo leaking from the cavity (i.e. unlocking of the cavity made the scattering disappeared)


Tomorrow Jenne and I will go into BS to tweak the alignment of the TEMP PRC flat mirror, and into ITMX to see if we can clean up the POP path.

  7948   Mon Jan 28 19:15:14 2013 ManasaUpdateScatteringScattering setup

 [Jan, Manasa]

We are trying to get some scattering measurements in the Y-arm cavity. We have removed one of  the viewport windows window covers of ETMY chamber and have installed cameras on a ring that clamps to the window. The window along with the ring attachment is covered with aluminium foil when not in use.

  7947   Mon Jan 28 19:07:45 2013 ManasaUpdateGeneralSN6 Laseroptik mirror - Tranmittance measurements

I repeated the measurements using NPRO instead of Crystalaser. I am attaching optical layouts for these measurements for future reference. 

Lesson learnt : Do not use Crystalaser for transmission measurements and always separate the transmitted main beam from other beams that result from the wedged surface of the mirror. 

trans_layout.png

Measurements match specs provided by Laseroptik

p-polarization

T percentage = 0.10% 42 deg
                            0.092% 44 deg
                            0.086% 46 deg
Minimum transmittance = 0.081% 52deg

s-polarization

T percentage = 0.048% 42 deg
                            0.047% 44 deg
                            0.047% 46 deg
Minimum transmittance = 0.047% 46 deg

sn6_trans0128.png
 

  7946   Mon Jan 28 17:59:02 2013 RijuUpdate Photodiode transimpedance

Summary: Measurement and plot of shot-noise-intercept-current for PDA255.

Motivation:It is to measure the shot noise intercept current for PDA255 - the MC transmission RF photodiode to get an idea for the noise current for the detector

Result: The final plot is attached here. The plot suggests that the value of shot-noise-intercept current is 3.06mA

Discussion:

The plot is for the measured data of Noise voltage (V/sqrt(Hz)) vs DCcurrent(A). The fitted plot to this measured data follows the noise equation

Vnoise = gdet* sqrt[ 2e (iDC+idet)] ,  where gdet= transimpedance of the PD in RF region as described in manual of PDA255 (i.e. 5e3 when it is not in High-impedance region).

On the other hand for DCcurrent calculation we must use the high-impedance value for the transimpedance i.e. 1e4 Ohm. idet is the shot noise intercept current.

For the rough calculation of the noise level we may use the following formulae:

Vnoise = gdet*sqrt[2e (iDC+idet)] = gdet*sqrt(2e in), when in=iDC+idet;

For say, in1=1mA; Vnoise1=gdet*sqrt(2e *in1)

and sqrt(2e *in1)~18pA/sqrt(Hz)

In current case dark noise is ~1.5e-7 V/sqrt(Hz)

Therefore dark current(in2) ~dark noise voltage/RF transimpedance = 30pA/sqrt(Hz)

i.e. sqrt(2e *in2)=30pA/sqrt(Hz)

i.e. sqrt(in2/in1)=30/18

therefore, in2~3mA (since in1=1mA)

For, iDC=0, in=idet.

Therefore the shot-noise-intercept current will be ~3mA

Then Vdc = in2*1e4 = 30V

According to the experiment  and also from the PDA255 manual the DC voltage level never goes beyond ~10V. Therefore following the photodiode characteristics(we work in reverse bias) we may infer that it can never become shot noise limited.

Also, from PDA255 manual, at 1650nm the dark noise is 30pW/sqrt(Hz) and the responsivity is 0.9A/W. Therefore the noise current level will be = noise power* responsivity ~27pA/sqrt(Hz). The value matches well with our expectation.

 

Attachment 1: shotnoiseinterceptpda255.pdf
shotnoiseinterceptpda255.pdf
  7945   Mon Jan 28 17:01:19 2013 DenUpdateLockingVideo of PRM-flat test cavity

What mode will you get if lock the cavity PRM - ITMY/ITMX/TEST MIRROR without PR2, PR3 and BS?

Is it possible to skip MC1, MC3 and lock the laser to this test cavity to make sure that this is not actuator/electronics noise?

  7944   Fri Jan 25 08:10:42 2013 SteveUpdateGeneralLaseroptik mirrors

 Here are the German plots. Unfortunately they are not logarithmic.

 Ed: Proprietary data removed. Use wiki (Koji)

Steve uploaded data to the 40m wiki  / Aux_Optics on 02-07-2013

 

 

  7943   Thu Jan 24 16:34:56 2013 ranaUpdateGeneralLaseroptik mirror - SN6

  I have two questions:

1) Are we sure that the T measurement is not being compromised by some systematic? i.e. some leakage is making the apparent T appear too high.

2) IF the T is really so high, how should we decide whether or not to use this one rather than the G&H? Is the 532 nm property more important than the high recycling gain?

  7942   Thu Jan 24 16:31:46 2013 SteveUpdatePEMbuilding exterier wall painted

The wood exteier walls, gutters and doors were painted at CES-Annex building #69

Attachment 1: IMG_1880.JPG
IMG_1880.JPG
  7941   Thu Jan 24 16:23:24 2013 KojiUpdateGeneralLaseroptik mirror - SN6

The mirror T is completely out of spec. We should find or request the data sheet of the mirror.

> We should consider buying a better rotation stage

I'm already on it

  7940   Thu Jan 24 15:16:50 2013 ManasaUpdateGeneralLaseroptik mirror - SN6

I repeated the transmittance measurements of Laseroptik SN6 at 1064nm. The rotation stage could only resolve 2 deg rotation (We should consider buying a better rotation stage).

s-polarization

Percentage transmittance

       0.177% 42 deg
       0.806% 44 deg
       0.57%   46 deg
       54.8%    0 deg

p-polarization

Percentage transmittance

       1.039% 42deg
       1.155% 44 deg
       1.159% 46 deg
        65.6%   0 deg

sn6_trans1.png 

  7939   Thu Jan 24 14:40:09 2013 SteveUpdateVACRGA backgroung

 RGA background with VM2 open to Maglev at day 37

 

Note: The PAN gauge of the  annulos is at atm.  Please do not vent this 200 ft long annulos line when you venting the annulos of a chamber. The chamber annulos should be closed off to this long 2"  OD. pipe before you vent the annulos of a chamber.

 

Attachment 1: Screenshot.png
Screenshot.png
  7938   Thu Jan 24 11:40:55 2013 JenneUpdateSUSETMX damping restored

Quote:

ETMX sus damping restored.  It is still noisy

 I should have elogged, but I turned off the watchdog to remind myself that iscex computer is still crashed.  "Turning on" the damping doesn't do anything since there aren't any signals going to the coils from the computer.

  7937   Thu Jan 24 11:31:45 2013 SteveUpdateSUSETMX damping restored

ETMX sus damping restored.  It is still noisy

  7936   Thu Jan 24 08:28:35 2013 SteveUpdateVACwhen to pumpdown ?

Reminder: I will be on vacation next week. We would have to put the access connector in tonight if I'm pumping down Friday, tomorrow.

  7935   Wed Jan 23 22:02:25 2013 ManasaUpdateGeneralLaseroptik mirror - SN6

Quote:

Got confused (even after I talked with Manasa).

The plot shows the number ~0.01 or less at 45deg. But the number is the text does not match with the plot.

Please use the logarithmic scale for the vertical axis.
And more points between 35 to 50 deg please (like ~1deg spacing)

Don't we have the data sheet from the coater? Can we request it?

I corrected the typo in the text...however, I agree the plot was lame...Will get the data sheet made tomorrow! 

  7934   Wed Jan 23 20:46:46 2013 Zen MasterUpdateLockingPR-flat cavity status - locks!

Quote:

I (with help from Q)

 Two quadratures working in harmony.

yinYang.png

  7933   Wed Jan 23 20:27:05 2013 RijuUpdate Photodiode transimpedance

Today I have repeated the expt for shot noise intercept current. Koji found that the Spectrum analyzer is going to saturation, so we have used one DC blocker (MCL - 15542 model) in PD signal.

I will analyze the data and report.

Ed by Koji: DC BLOCK is  BLK-89-S

  7932   Wed Jan 23 20:24:05 2013 KojiUpdateGeneralLaseroptik mirror - SN6

Got confused (even after I talked with Manasa).

The plot shows the number ~0.01 or less at 45deg. But the number is the text does not match with the plot.

Please use the logarithmic scale for the vertical axis.
And more points between 35 to 50 deg please (like ~1deg spacing)

Don't we have the data sheet from the coater? Can we request it?

  7931   Wed Jan 23 19:05:16 2013 JenneUpdateLockingPR-flat cavity status - locks!

Status update

I (with help from Q) have redone the POP path on the ITMX table.  1" iris is a little too small, so I took it out.  2" lens moved to be centered on POP beam.  2" Y1 didn't need moving.  Straight refl from the 2" Y1 was aligned on to a PDA10CS (set to 70dB). This PD is blocking the usual POP55 diode.  BS which sends beam to camera was moved to allow room for the new temp DC PD.  Refl from this BS goes to the POP camera, which was moved so that the POP beam takes up most of the camera.  BS that would normally take half of the camera's beam and send it to POP22 (Thorlabs PD) is removed, so no beam to POP22.

Also, I have taken the output of the PDA10CS and hijacked the "POP110" heliax cable.  This was connected to this Thorlabs PD which is used as POP22.  (Kiwamu and I had long-term borrowed the 110 demod board for an AS 110 diode, so the "POP110" heliax was really only serving POP22.) There are yellow labels on the new temp and old regular cables, so we can undo my hack.  Similarly, on the other end of the heliax at the LSC rack, I have taken the heliax's output and sent it to the POPDC input on the whitening board.  Thus, the regular POPDC SMA cable is unplugged, but labeled again with big yellow labels.

In other news - the PR-flat cavity locks!!! 

Koji and I coarsely rotated the REFL11 phase such that the signal is predominantly in the I phase.  We set the LSC input matrix to use REFL11I for PRCL, and the output matrix is set to actuate on PRM.  Then we set the gain to -0.005, and it locked!!!!

 

EDIT:  I turned back on the PRM oplev (after Manasa aligned it and redid the out-of-vac oplev layout a bit), and the motion of the cavity is slightly reduced, although there's still a lot going on.  The cavity is vaguely well aligned, although it's time to go make sure that the beams are still on the REFL and TRANS PDs.  However, it's dinner time.

  7930   Wed Jan 23 18:16:11 2013 ManasaUpdateGeneralLaseroptik mirror - SN6

 I repeated the transmittance measurements of LaserOptik SN6 @1064nm.

Transmittance for s-polarization 

0 deg - 0.524
45 deg - 0.055

Transmittance for p-polarization

0 deg - 0.515
45 deg - 0.1047 0.01047

Raji's measurements are here.

Attachment 1: sn6_trans.png
sn6_trans.png
  7929   Wed Jan 23 11:43:19 2013 RijuUpdate Photodiode transimpedance

Quote:

- The data should be plotted in a log-log scale.
- The data points were only taken in the high current region.

- The plot may suggest that the amplifier saturate at the RF.

PDA255 has the nomial transimpedance gain of 10^4 Ohm.
The DC current of 10^-3 gives the output of 10V.
This plot may tell that the saturation starts even at the 1/10 of the full DC range.

The plot doesn't have many points below 0.1mA.
Consult with my plots for the similar measurements.
The measured points are logarithmically spaced. Use the same technique.

- It is also very unknown that how the noise level is calculated. No info is supplied in the plot or the elogs.

 Here I am attaching the plot in loglog scale. I have taken the data-points from no light condition to the maximum light condition, the minimum variation possible in the current supply was 0.01A. The noise was visibly decreasing at higher light level.

For the noise level calculation I took the average of total noise in the range 7-60MHz. For each range the formula used was

noisevalue= sqrt(data(:,2)*100)/sqrt(2)/sqrt(channel BW);     -- this conversion is needed since the data was collected in the 2 column format: frequency, spectrum(W).

Attachment 1: pda255shotnoiseintercept1.pdf
pda255shotnoiseintercept1.pdf
  7928   Tue Jan 22 19:53:01 2013 JenneUpdateLockingPR-flat cavity status - not locked

The PR-flat cavity is flashing, although not locked.  I am too hungry to continue right now.

I put the FI_Back camera on a tripod, looking at the back of the Faraday.  The beam that Jamie and I were working with on Friday was clipped going back through the Faraday.  I twiddled the TT2 and PRM pointing such that the beam is retroreflecting, and getting back through the Faraday, and the cavity is still flashing.  I then redid the REFL path on the AS table a little bit.  The beam is currently going to the REFL camera, as well as REFL11 and REFL55. 

Some notes about the AS table:  The Y1 separating the main REFL beam from the REFL camera beam was mounted 90 degrees (rotated about the beam's axis) from what it should be.  I fixed it, so that the straight-through beam that goes to the camera is not clipped by the edge of the mount.  The reason (I think) this mirror was mounted backwards is that when mounted correctly, the back of the mount and the knobs interfere with the AS beam path.  I solved this by rotating the first out-of-vac REFL mirror a small amount so that the REFL and AS beams are slightly more separated. 

I am not seeing any nice PDH signal on dataviewer, so I went to check the signal path for the PDs.  The 11MHz marconi is on and providing RF, the EOM is plugged in to 11, 55 and 29.5 signals (no aux cavity scan cables are plugged in).  Both of the RF Alberto boxes are on.  I measured the RF output of both REFL11 and REFL55, although after the fact I realized that I was BAD, and had not found a 'scope that lets me change the input impedance to 50 ohms.  BAD grad student.  However, since I have numbers, I will post them, despite their being not quite correct:

284mVpp at 11MHz out of REFL11.  This is -6.9dBm

2mVpp at 55MHz out of REFL55, measured by 'scope

So, I can clearly see the 11MHz on the 'scope, and can see a very noisy, small 55MHz signal on the 'scope.  I need to think over dinner about what level of signal we should be sending to the demod boards, and whether or not I need more power coming out of the RFPDs.  There is a wave plate and PBS before beam goes to any of the REFL PDs, presumably to ensure that none of them get fried when we're at high power.  If I need more signal, I suspect I can rotate the wave plate and let more light go to the diodes.

  7927   Tue Jan 22 19:51:52 2013 KojiUpdate Photodiode transimpedance

- The data should be plotted in a log-log scale.
- The data points were only taken in the high current region.

- The plot may suggest that the amplifier saturate at the RF.

PDA255 has the nomial transimpedance gain of 10^4 Ohm.
The DC current of 10^-3 gives the output of 10V.
This plot may tell that the saturation starts even at the 1/10 of the full DC range.

The plot doesn't have many points below 0.1mA.
Consult with my plots for the similar measurements.
The measured points are logarithmically spaced. Use the same technique.

- It is also very unknown that how the noise level is calculated. No info is supplied in the plot or the elogs.

  7926   Tue Jan 22 17:29:29 2013 RijuUpdate Photodiode transimpedance

Riju

Summary:  I am stuck with the measurement of shot-noise-intercept-current of PDA255. Seeking help.

Motivation: It is to measure the shot noise intercept current for PDA255 - the MC transmission RF photodiode to get an idea for the noise current for the detector.

Method: It is as described in the elog  7907 

Result: The plot is attached here.

Discussion: The result I got is really unexpected, the noise voltage should increase with the DC current level that corresponds to the increment of light level too. But actually it is decreasing. Three times I have repeated this experiment and got the same result. I want some suggestion on this regard.

Attachment 1: pda255shotnoiseintercept.pdf
pda255shotnoiseintercept.pdf
  7925   Tue Jan 22 11:09:58 2013 SteveUpdateTip - TiltSS mirror holder ordered

Koji's design of the SS  2" mirror holder with flexure spring optic retainer  like Polaris-K1 has been ordered. We are getting just one to see it's effect on the hysteresis.

Attachment 1: D1201574_v1_Tip-Tilt_Mirror_Holder.SLDDRW
  7924   Tue Jan 22 09:18:50 2013 SteveUpdateTip - Tilttip tilt bases ordered

The corrected drawing base for tip tilt with coils are going to the shop. The will be back by the end of the week.

Attachment 1: TTbc.PDF
TTbc.PDF
  7923   Tue Jan 22 09:10:19 2013 SteveUpdatePEMbig foot is dirty

Please wipe, clean car wheels and wear booties entering the 40m lab.

Obviously this person has no idea about our clean room rules.

Attachment 1: IMG_1873.JPG
IMG_1873.JPG
Attachment 2: IMG_1874.JPG
IMG_1874.JPG
  7922   Sat Jan 19 18:23:31 2013 ranaUpdateComputersAll front ends but c1lsc are down

After sshing into several machines and doing 'sudo shutdown -r now', some of them came back and ran their processes.

After hitting the reset button on the RFM switch, their diagnostic lights came back. After restarting the Dolphin task on fb:

"sudo /etc/init.d/dis_networkmgr restart"

the Dolphin diagnostic lights came up green on the FE status screen.

iscex still wouldn't come up. The awgtpman tasks on there keep trying to start but then stop due to not finding ADCs.

 

Then power cycled the IO Chassis for EX and then awtpman log files changed, but still no green lights. Then tried a soft reboot on fb and now its not booting correctly.

Hardware lights are on, but I can't telnet into it. Tried power cycling it once or twice, but no luck.

Probably Jamie will have to hook up a keyboard and monitor to it, to find out why its not booting.

FE.png

P.S. The snapshot scripts in the yellow button don't work and the MEDM screen itself is missing the time/date string on the top.

  7921   Sat Jan 19 16:02:28 2013 ranaOmnistructureElectronicsPS cleanup

During our 'Women in Physics' tours today, we were reminded that there are several bench power supplies being used as permanent inside.

Some are being used to power PZTs, AOMs, VCOs, RFPDs, etc. On Wednesday, after the meeting, we will all go inside and remove one and replace it with a fused, professional wiring to the rack power supplies. The temporary ones must be removed.

  7920   Sat Jan 19 15:05:37 2013 JenneUpdateComputersAll front ends but c1lsc are down

Message I get from dmesg of c1sus's IOP:

[   44.372986] c1x02: Triggered the ADC
[   68.200063] c1x02: Channel Hopping Detected on one or more ADC modules !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
[   68.200064] c1x02: Check GDSTP screen ADC status bits to id affected ADC modules
[   68.200065] c1x02: Code is exiting ..............
[   68.200066] c1x02: exiting from fe_code()

Right now, c1x02's max cpu indicator reads 73,000 micro seconds.  c1x05 is 4,300usec, and c1x01 seems totally fine, except that it has the 02xbad.

c1x02 has 0xbad (not 0x2bad).  All other models on c1sus, c1ioo, c1iscex and c1iscey all have 0x2bad.

Also, no models on those computers have 'heartbeats'.

C1x02 has "NO SYNC", but all other IOPs are fine.

I've tried rebooting c1sus, restarting the daqd process on fb, all to no avail.  I can ssh / ping all of the computers, but not get the models running.  Restarting the models also doesn't help.

upside_down_cat-t2.jpg

c1iscex's IOP dmesg:

[   38.626001] c1x01: Triggered the ADC
[   39.626001] c1x01: timeout 0 1000000
[   39.626001] c1x01: exiting from fe_code()

c1ioo's IOP has the same ADC channel hopping error as c1sus'.

 

  7919   Fri Jan 18 15:08:13 2013 jamieUpdateAlignmentalignment of temporary half PRC

[jenne, jamie]

Jenne and I got the half PRC flashing.  We could see flashes in the PRM and PR2 face cameras.

We took out the mirror in the REFL path on the AP that diverts the beam to the REFL RF pds so that we could get more light on the REFL camera.  Added an ND filter to the REFL camera so as not to saturate.

  7918   Fri Jan 18 12:08:08 2013 KojiUpdateLockingDust?

No

Quote:

My question is, could a few scattering centers cause the behavior that we are seeing?

 

  7917   Fri Jan 18 09:54:18 2013 JenneUpdateLockingPRM - Flat mirror cavity

Quote:

I do, however, see flashes on the face of the PRM.  Movie saved, will post soonly.

 Dang it.  I didn't confirm that the movie was good, just that it was there.  It's corrupted or something, and won't play.  I'll just have to make a new movie today after I realign the cavity.

  7916   Fri Jan 18 00:41:34 2013 JenneUpdateLockingDust?

I was thinking tonight about more possible reasons that our PRC sucks, and I wonder if dust on the BS could create the problem.

Historically, Kiwamu and I found a few dust particle scattering centers every time we inspected the test masses before drag wiping. Sometimes, there would be one frustratingly close to the center of the optic. I'm not sure if we ever made note of how many we saw and where they were, except out loud to the assembled crowd.

Anyhow, the BS is the only IFO optic that was not replaced, so I'm not sure how long it has been since it was cleaned. If the PR-flat cavity looks okay and we take out the BS to do a PRM-ITMY cavity, we should inspect the beam splitter.

Also, the PRM could need cleaning, but at least it has been drag wiped within recent memory.

My question is, could a few scattering centers cause the behavior that we are seeing?

 

EDIT:  List o' elogs....

 

Elog 5301 - Some details on dust seen on ITMs and ETMs, Aug 2011.

Elog 4084 - Kiwamu's in-situ drag wiping how-to, with details on some of the dust we saw. Dec 2010.

Elog 3736 - PRM drag wiped before suspension (Oct 2010)

Elog 3111 - June 2010, BS drag wiped.

  7915   Thu Jan 17 19:33:53 2013 ManasaUpdateAlignmentPRM oplev

 We had to work on redesigning the oplev layout in BSC when I found that the positions of the mirrors were clipping IPPOS and the green beam while updating the CAD layout.  

To avoid any clipping, the prm oplev beam is steered into the vacuum by an oplev mirror and out of vacuum through 3 steering mirrors. The table weights had to be moved to allow room for the oplev mirrors. Hence table had to be re-leveled. I will update the CAD drawing with the current position of the mirrors and will reconfirm that the new mirrors are not in the way of any of the beams. In-vac photos are updated in picasa.

  7914   Thu Jan 17 17:04:49 2013 SteveUpdateVACvac monitor screen changed

The COVAC_MONITOR.adl was changed. Ion pump labeled as disconnected means: ion pump controller power turned off and ion pump gate valve control connection disconnected.

RP2 roughing pump labeled disconnected means:  hardware  disconnected.

The actual operational, valve control screen has not been changed yet.

Attachment 1: screenmod.png
screenmod.png
  7913   Thu Jan 17 15:48:21 2013 JenneUpdateLockingPRM - Flat mirror cavity

 

 2" G&H mirror is installed on a DLC mount just in front of the BS.  I had to remove one of the 4 BS dog clamps, so we must put it back when we are finished with this test.

I aligned the G&H mirror such that the reflected beam is overlapped with the incident beam, and I aligned the PRM such that the regular REFL beam is retro-reflected.  This is the same as getting the beam bouncing off the PRM back to the G&H to be overlapped.

I then saw flashes of the cavity, when I held a card with a hole in the cavity, so the beam was going through a small aperture in the card, but I still saw flashes.  I was not able to see flashes on the IR card transmitted through the G&H mirror.

I also cannot see any flashes or scattered light on the face of PR2 camera.

I do, however, see flashes on the face of the PRM.  Movie saved, will post soonly.

Light is coming out of REFL on the AS table, but it's clipped somewhere....needs investigation/work before we can lock.

I also didn't see anything at the POP port with a card, but I'm hopeful that perhaps with a camera I'll see something.

  7912   Thu Jan 17 11:01:19 2013 JenneUpdateAlignmentYesterday's alignment work

[Jamie, Jenne, Manasa]

Yesterday's goal was to get the input beam centered on the PRM, the BS and ETMY simultaneously. 

Steve helped us remove the ETMY door first thing in the morning.  We then iterated with TT1, MMT1 and TT2 to try to get the beam centered on all the optics.  We were using MMT1 instead of TT1 for a while, so that we could keep TT1 in the center of its range, so that we had more range to use once we pump down.  Also, at one point, the beam was high on PRM, centered on BS, and high on ETMY, so Jamie poked PR3 a little bit.  This helped, although we closed up for lunch / group meeting soon after, so we didn't finalize any alignment stuff.

We decided to leave the rest of the full IFO alignment alone until after the PRM-flat test.

  7911   Thu Jan 17 01:27:54 2013 Tara(?)SummaryIOONoise budget for MC

Quote:

I missed the point.
Do you mean that we can measure the coating thermal noise of the ref cav at the 40m as the IMC is quiet enough?

 Yes, it should be. However, what I did was calculating thermal noise of MC. I'm not sure about the 40m IMC's actual noise level. The plot in the entry was taken from LLO's MC in 2003.

  7910   Thu Jan 17 00:17:31 2013 JenneUpdateLockingExpected PRC gains

Quote:

Why would we use such a bad optic in our recycling cavity? Is 1.5% the spec for these mirrors? Is this the requirement that Kiwamu calculated somehow? Did anyone confirm this measurement?

I can't believe that we'll have low noise performance in a RC where we dump so much power.

 Yeah, Koji mentioned in response to Raji's measurements several months ago that the LaserOptic mirros were pretty far out of spec. We should probably redo the measurement to confirm.

  7909   Wed Jan 16 20:27:16 2013 ranaUpdateLockingExpected PRC gains

Why would we use such a bad optic in our recycling cavity? Is 1.5% the spec for these mirrors? Is this the requirement that Kiwamu calculated somehow? Did anyone confirm this measurement?

I can't believe that we'll have low noise performance in a RC where we dump so much power.

  7908   Wed Jan 16 19:08:51 2013 KojiSummaryIOONoise budget for MC

I missed the point.
Do you mean that we can measure the coating thermal noise of the ref cav at the 40m as the IMC is quiet enough?

  7907   Wed Jan 16 18:58:08 2013 RijuUpdate Photodiode transimpedance

Today I have taken the reading for shot noise intercept current for the PDA255 - MC transmission RF PD. To do that I have put an incandescent bulb (JKL lamps, 222 bulbs, voltage and current rating 2.25V and 0.25A) in front of the PD and varied the current through it from 0A to 0.29A at 2.2V. I measured the corresponding DC voltage and took the noise data (4395A spectrum analyzer/ format noise, channel attenuation 0dB) through GPIB .

I will process the data and upload the result soon.

  7906   Wed Jan 16 18:52:49 2013 JenneUpdateLockingPRM - Flat mirror cavity plan

Game plan:

* Put 2" G&H mirror into BS chamber, in front of BS.

* Align it, lock cavity using an existing REFL PD.

* Align POP setup so I can use POP camera to take image of transmitted cavity mode, and actually take that image.

* Take image of face of PR2.

* Measure finesse of cavity using POP, or a Thorlabs PD at POP (looking at transmission through PR2) by scanning PRM, and infer cavity gain....compare with values in elog 7905.

* If time / inclination allow, take beam scan measurements of the REFL port.

I will not be able to do as was done in elog 6421 to look at the beam size at POP for non-resonating beams.  I expect ~0.1uW of light at POP in the non-resonant case:  100mW * 5.5% * 20ppm = 0.11microwatts.

 

  7905   Wed Jan 16 18:08:06 2013 JenneUpdateLockingExpected PRC gains

I was calculating the power recycling gains we expect for different versions of the PRC, and I am a little concerned that we aren't going to have much gain with the new LaserOptik mirrors.

I'm using

                       t_PRM^2

G =  -------------------------------------------

       (1 - r_PRM * r_PR2 * r_PR3 * r_end)^2

 

from eqn 11.20 in Siegman.

r_end is either the ITM (for a symmetric Michelson) or the flat mirror that we'll put in (for the PR-flat test case).

r = sqrt( R ) = sqrt( 1 - T ) for mirrors whose power transmission is the quoted value.

 

Some values: 

t_PRM^2 = T_PRM = 0.055   --------->   r_PRM = sqrt( 1 - 0.055 )

T_G&H = 20e-6   ---->   r_G&H = sqrt( 1 - 20e-6 )

T_LaserOptic = 0.015 (see elog 7624 where Raji measured this...1.5% was the best that she measured for P polarization.  Elog 7644 has more data, with 3.1% for 40deg AoI) -------> r_LasOpt = sqrt( 1 - 0.015 ) or sqrt( 1 - 0.031)

T_ITM = 0.014 -----------> r_ITM = sqrt( 1 - 0.014 )

 

Some calculations with 1.5% LaserOptik transmission:

G_PRC_2G&H = 45

G_PRC_G&H_LasOpt = 31

G_PRM_flatG&H = 51

With the 3% LaserOptik transmission:

G_PRC_G&H_LasOpt = 22

G_PRM_flatG&H = 30

More ideal case of just PRM, flat mirror (either ITM or G&H), ignoring the folding mirrors:

G_PRM_ITM = 45

G_PRM_flatG&H = 70

 

Punchline:

If the LaserOptik mirror has 1.5% transmission at ~45 degrees, the regular PRC expected gain goes down to 31, from 45 with both folding mirrors as G&Hs.

  7904   Wed Jan 16 10:57:37 2013 taraSummaryIOONoise budget for MC

I calculated thermal noise in mode cleaner (MC) mirrors and compared it with the measured MC noise. Thermal noise won't be a significant noise source for MC.

== Motivation==

 There is an idea of using MC and a refcav to measure coating thermal noise. One laser is frequency locked to MC, another laser is locked to an 8" refcav. Then the two transmitted beams are recombined so that we can readout the frequency noise. In this case, the transmitted beam from MC is a better reference (less frequency noise) than the beam from refcav. However, we need to make sure that we understand the noise sources, for example brownian noise, thermoelastic noise in both substrates and coatings, in MC more thoroughly.

==Calculation==

I used Rana's code for MC's technical noise sources from, svn. The same plot can be found in appendix C of his thesis. Then I added my calculation to the plot.  Jenne pointed me to 40m:2984 for the spot size and the cavity length. The spot radius on MC1 and MC3 is ~ 1.5mm, and ~3.4 mm@MC2, The round trip length is ~27m, thus the frequency fluctuation due to thermal noise is lower than that of refcav by 2-3 orders of magnitude. I calculated Brownian noise in coatings, Brownian noise in substrate, Thermoelastic noise in substrate. I assumed that the coatings are SiO2/Ta2O5, quarter stacks, coatings thickness for MC1/3 = 5um, for MC2 = 8um. The code can be found in the attachment.

mc_nb_TN.png

==result==

Total thermal noise on MC (Brownian + Thermoelastic on substrate and coatings of MC1-MC3) is plotted in dashed red. It is already below 10^-5 Hz/rtHz at ~20 Hz. This is sufficiently low compared to other noise sources. Beat signal from CTN measurement with 8" cavities is plotted in pink, the estimated coating brownian noise is plotted in a yellow strip. They are well above the measured MC noise between 100 Hz to a few kHz. Measuring coating thermal noise on 8" refcav seems plausible with this method. We can beat the two transmitted beams from IMC and refcav and readout the beat signal to extract the displacement noise of refcav. I'll discuss this with Koji if this is a good surf project.

 mc_nb_TN_2013_01_18.png

[the internal thermal noise in the original plotted is removed and replaced with the total thermal noise plot instead]

 note:I'm not sure about the current 40m MC configuration. The parameters used in this calculation are summarized in mcnoiseS2L1.m (in the svn page).

 

Attachment 2: mc_nb_TN.png
mc_nb_TN.png
Attachment 3: mc_nb_TN.fig
Attachment 4: MC_nb.m.zip
  7903   Wed Jan 16 08:29:45 2013 SteveUpdateSUSPRM damping restored

PRM oplev servo turned off.  OLPIT servo gain 0.15 and OLYAW  -0.3 set to ZERO.  PRM damping restored

  7902   Tue Jan 15 20:00:42 2013 ManasaUpdateAlignmentAdjustment of active TTs and input alignment

 

Just for reference! The changes made to the TT matrix in order to fix the polarity problem:

The old matrix values are mentioned in elog!

 

PIT    YAW                   New                   

Pit slider           |  -100   -100  |  UL  

     0               | -100    100   |  LL

Yaw slider           |  100   -100   |  UR

    0                |  100    100   |  LR

 

 

 

 

 

 

 

  7901   Tue Jan 15 19:26:35 2013 jamieUpdateAlignmentAdjustment of active TTs and input alignment

[Jamie, Manasa, Jenne]

We started by verifying that the tip-tilts were getting the correct signals at the correct coils, and were hanging properly without touching.

We started with TT2.  It was not hanging freely.  One of the coils was in much further than the others, and the mirror frame was basically sitting on the back side yaw dampers.  I backed out the coil to match the others, and backed off all of the dampers, both in back and the corner dampers on the front.

Once the mirror was freely suspended, we borrowed the BS oplev to verify that the mirror was hanging vertically.  I adjusted the adjustment screw on the bottom of the frame to make it level.  Once that was done, we verified our EPICS control.  We finally figured out that some of the coils have polarity flipped relative to the others, which is why we were seeing pitch as yaw and vice-versa.  At that point we were satisfied with how TT2 was hanging, and went back to TT1.

Given how hard it is to look at TT1, I just made sure all the dampers were backed out and touched the mirror frame to verify that it was freely swinging.  I leveled TT1 with the lower frame adjustment screw by looking at the spot position on MMT1.  Once it was level, we adjusted the EPICS biases in yaw to get it centered in yaw on MMT1.

I then adjusted the screws on MMT1 to get the beam centered at MMT2, and did the same at MMT2 to get the beam centered vertically at TT2.

I put the target at PRM and the double target at BS.  I loosened TT2 from it's base so that I could push it around a bit.  Once I had it in a reasonable position, with a beam coming out at PR3, I adjusted MMT1 to get the beam centered through the PRM target.  I went back and checked that we were still centered at MMT1.  We then adjusted the pitch and yaw of TT2 to get the transmitted beam through the BS targets as clear as possible.

At this point we stopped and closed up.  Tomorrow first thing AM we'll get our beams at the ETMs, try to finalize the input alignment, and see if we can do some in-air locking.

The plan is still to close up at the end of the week.

  7900   Tue Jan 15 01:41:40 2013 ranaUpdateSUSTT

 That seems like easily enough range; as long as we can put the TT into the middle of their range to start with we should be OK.

We should consider instrumenting the leakage transmission through all TT with a bare QPD on a stick. We can then use those sensors to monitor the spot positions within the input mirrors as well as the PRC / SRC.

ELOG V3.1.3-